OpenStudy (anonymous):

next numbers are.. ? 4, 12, 36, 108, . . . A. 304, 912, 2,736 B. 324, 972, 2,916 C. 354, 1,062, 3,186 D. 332, 996, 2,988

OpenStudy (anonymous):

a(n)=(n-1)*3 108*3=324 324*3=972 972*3=2,916 (B)